How do i work out a problem when given
How long will it take for the population of a certain country to double
if its annual growth rate is 5 percent?
First I plugged in 5 (for 5%) Log in my calculatorand the answer was 0.l698970004.
Do i move the decimal over 2 places to give me 16.989 years or
rather that it would take 17 years for the population to double.
I have researched everthing and am unable to come up with any samples on how to work out any problems given just the percent without a population number? Can you please help me understand how to do this?
Is it just a calculator problem?
Any help would be greatly appreciated.
Thank You.

Answer by checkley77(12569)

i=p(1+r)^t
in this equation (1+r)^t must=2 to double
the principle.
(1+.05)^t=2
1.05^t=2
[here the 't' value was obtained by the trial & error methoid as my calculator doesn't have the proper function.]
t=14.207 or it wil take 14.207 years to double your investment @ 5% annually.